How would I solve this?
An investment of $1500 paid interest of $45 during a certain period. How much interest would an investment of $2000 invested at the same rate for the same length of time pay?

Answer by stanbon(75887)

How would I solve this?
An investment of $1500 paid interest of $45 during a certain period. How much interest would an investment of $2000 invested at the same rate for the same length of time pay?
--------------
I = Prt
---
45 = 1500rt
I = 2000rt
-------
Set up a proportion:
I/45 = 2000/1500
I = 45(2000/1500)
I = 45(4/3)
---------------------
Interest on $2000 = $60
=============================
Cheers,
Stan H.
=============